2) find the valeue of k if |K 3|
|K 2K|=0
1
Expert's answer
2013-01-29T10:49:12-0500
Determinant of that matrix is det = K*2K - 3K= K^2-3K = (K-3) K
So
(K-3) K = 0whence K=3 or K=0.